I don't know how to set the equations up. I tried using the box method. D=RT but I keep coming up with different numbers everytime.
Mitzi Fulwood traveled for 2 hr at a constant speed. Because of roadwork, she reduced her speed by 7 mph for the next 2 hr. If she traveled 206 mi, what was her speed on the first part of the trip?

Answer by checkley77(12844)

2x+2(x-7)=206
2x+2x-14=206
4x=206+14
4x=220
x=220/4
x=55 mph for the first 2 hours.
55-7=48 mph for the next 2 hours.
proof
2*55+2*48=206
110+96=206
206=206